OpenStudy (anonymous):

How would you change the base formula to rewrite a expression using common logarithms? log(base7)8

OpenStudy (tkhunny):

Just do it? \(\log_{7}(8) = \dfrac{\log(8)}{\log(7)}\)

OpenStudy (jdoe0001):

\(\bf \textit{log change of base}\implies log_ab=\cfrac{log_{\color{red}{ c}}b}{log_{\color{red}{ c}}a}\)

OpenStudy (anonymous):

So is the base always on the bottom of the fraction

OpenStudy (jdoe0001):

yes and "c" could be any number really, so long is the same above and below

OpenStudy (anonymous):

ahh okay, i get it. But can i ask one more question?

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

ok 1 second

OpenStudy (jdoe0001):

1

OpenStudy (jdoe0001):

hehe

OpenStudy (anonymous):

lol, how would you write this as a single logarithm. log\[\log_{5}4 +\log_{5}3 \]

OpenStudy (jdoe0001):

http://www.chilimath.com/algebra/advanced/log/images/rules%20of%20exponents.gif <-- notice the 1st rule there

OpenStudy (anonymous):

Oh ok so would it be \[\log_{5} (4x3) \] ?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

:D ahh alright thank you you!

OpenStudy (jdoe0001):

yw